OpenStudy (anonymous):

in the united states 154000000 tons of garbage are produced annually. on an averge about how many pounds is that each month for each person in the us? the population of the us is about 250 million

OpenStudy (anonymous):

thank you ash

OpenStudy (ash2326):

Total Garbage= 154000000 Tonnes annually In a month = 154000000/12= 12833333.3 tonnes Population is US= 25 million or 250000000 So average garbage per month per person= 12833333.3/250000000= 0.0513 tonnes

OpenStudy (anonymous):

wow!

OpenStudy (anonymous):

thx so much

OpenStudy (ash2326):

welcome ori:)

OpenStudy (anonymous):

:)